OpenStudy (anonymous):

derrivative of g(x)= (2x+4)/(3x-1)

OpenStudy (anonymous):

You have to use the product rule. Do you know this rule?

OpenStudy (anonymous):

kinda. when i calculated it, it didnt turn out right :/

OpenStudy (anonymous):

Product Rule: g(x)=vu’+uv'

OpenStudy (anonymous):

Actually my mistake it is quotient rule.

OpenStudy (anonymous):

g(x)=(vu’-uv’)/v^2

OpenStudy (anonymous):

So in this case g(x)=(2x+4)/(3x-1) So v=3x-1 u=2x+4 v’=3 u’=2 Can you calculate now?

OpenStudy (anonymous):

yeah. thanks